OpenStudy (anonymous):

Express the domain of the following function in interval notation. (sqrt) 3-4x

OpenStudy (anonymous):

@welshfella can you please help

OpenStudy (welshfella):

Hint: because of the square root 3 - 4x cannot have a negative value

OpenStudy (anonymous):

so x has to be a negative number?

OpenStudy (welshfella):

no 3 - 4x >= 0 solve this inequality to find valid values of x This will be the domain

OpenStudy (anonymous):

so, x\[x \ge-3/4\]

OpenStudy (welshfella):

no you have gone wrong somewhere 3 - 4x >= 0 -4x >= -3 did you get that?

OpenStudy (anonymous):

yes i get it

OpenStudy (anonymous):

so its positive 3/4

OpenStudy (welshfella):

right now you divide both sides by -4 . But dividing by a negative flips the inequality sign around

OpenStudy (anonymous):

so \[x\]

OpenStudy (welshfella):

yes but is it >= 3/4 or <= 3/4?

OpenStudy (anonymous):

<=

OpenStudy (welshfella):

right now you have to write that in interval form

OpenStudy (anonymous):

okay one moment

OpenStudy (anonymous):

\[(-\infty, 3/4)\]

OpenStudy (anonymous):

]

OpenStudy (welshfella):

almost right since 3/4 is included in the domain the last parentheses should be ]

OpenStudy (anonymous):

alright got it

OpenStudy (anonymous):

thank you!

OpenStudy (welshfella):

yw